Crew 236 – EVA Report – 24-Dec-2021

[Commander's EVA Report – December 24th]

Crew 236 EVA Report 24-DEC-2021

EVA # 5

Author: Kasey Hilton

Purpose of EVA: Evaluation of soil and ground conditions for additional potential habitat locations

Start time: 13:00

End time: 16:00

Narrative: EVA was canceled this afternoon due to winds and rain in the morning, making the ground very muddy. Will try this EVA again on 25DEC2021.

Destination: End of Galileo Rd/Area F

Coordinates (use UTM NAD27 CONUS): N4251500, E519750

Participants: Kasey (CMD), Vladimir (GHO), Dylan (ASTRO)

Road(s) and routes per MDRS Map: Cow Dung Rd north to Galileo Rd, then heading down into area F around Compass Rock

Mode of travel: Driving and walking

Crew 236 Operations Report 24-Dec-2021

[Operations Report – December 24th]

Crew 236 Operations Report 24-DEC-2021

SOL: 5

Name of person filing report: Pavithra Ravi

Non-nominal systems:

1. Suits and radios — the suit issue reported on yesterday’s report persists.

Notes on non-nominal systems: The backpack was opened and the voltage
across the battery leads was checked with a multimeter, yielding 11.48
V. Since this is not too far from the ideal value of 12 V, the issue
may not be with the battery. When the air was turned on to max, one of
the fans did not turn at all, while the second fan spun very slowly.
Nothing appeared to be burned or abnormal.

2. Toilet — the leaking issue persists when the water valve isn’t turned off.
Today, we attempted to fix the leak by reapplying Teflon tape to the
threading on the fittings near the leak spot. Then, we proceeded to
tighten the fittings once more and then we turned on the water valve.
After about an hour, the leak started once more, though now it is
dripping from a different spot. While the water was initially coming
out just behind the black fitting (shown in with the blue arrow in the
attached photo), now it is coming out where the black arrow is.

ROVERS

Spirit rover used: No

Hours: (before EVA): N/A

Beginning charge: (Before EVA) N/A

Ending charge: (On return from EVA, before recharging) N/A

Currently charging: Yes

Opportunity rover used: No

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: Yes

Curiosity rover used: No

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Perseverance rover used: No

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: Yes

General notes and comments: None

Summary of Hab operations: Toilet — the leaking issue persists when
the water valve isn’t turned off.
Today, we attempted to fix the leak by reapplying Teflon tape to the
threading on the fittings near the leak spot. Then, we proceeded to
tighten the fittings once more and then we turned on the water valve.
After about an hour, the leak started once more, though now it is
dripping from a different spot. While the water was initially coming
out just behind the black fitting (shown in with the blue arrow in the
attached photo), now it is coming out where the black arrow is.

WATER USE:

Water (static tank): approx. 334 gallons

Water (loft tank): 50 gallons

Water Meter: 01540146 units

Static to Loft Pump used – yes or no: Yes. It was filled from 19 to 55
gallons at about 17:30 today.

Static tank pipe heater (on or off): Off

Static tank heater (on or off) Off

Toilet tank emptied (no or yes): No. The tank was last emptied at
about 20:45 on Dec 22nd. This was done after the report on the 22nd
was sent and accidentally left out on yesterday’s report.

Summary of internet: Nominal

Summary of suits and radios: The suit issue reported on yesterday’s
report persists. The backpack was opened and the voltage across the
battery leads was checked with a multimeter, yielding 11.48 V. Since
this is not too far from the ideal value of 12 V, the issue may not be
with the battery. When the air was turned on to max, one of the fans
did not turn at all, while the second fan spun very slowly. Nothing
appeared to be burned or abnormal.

Summary of GreenHab operations: Nominal

WATER USE: 9.5 gallons were used today. This leaves about 203.5
gallons in the tank.

Heater (on or off): On

Supplemental light (on or off): On

Harvest: 2g basil, 37g tomato, 1g green onion

Summary of ScienceDome operations: Nominal

Dual split, Heat or AC (on or off): off

Summary of RAM operations: Tools used, work done

A multimeter was used to obtain a voltage reading for the suit that
was malfunctioning. Teflon tape and a wrench were used to fix the
toilet leak.

Summary of any observatory issues: Nothing new to report.

Summary of health and safety issues: Health status nominal. A minor
incident occurred which has been detailed in the HSO report.

Questions, concerns and requests to Mission Support:

There is trash left in the back airlock.

We require additional support on solving the suit and toilet issues.

Journalist Report – December 24th

Journalist Report
Ben Durkee, Crew 236 Journalist

Sol 05

We’ve begun to develop our own Martian dialect. It sounds a lot like our collective worst impressions of a cockney British accent. I’m genuinely shocked that it hasn’t become annoying yet – I think the space madness has begun to set in and it manifests in the most mysterious ways. If the Health & Safety Officer had to write psychological reports, he would have A LOT of material at this point.

On the flip side, today was a relatively slow day on Mars which was a boon to our personal projects (and sleep schedules), but doesn’t give me a lot to work with. C’mon, Ben, think! Um… lovely weather we’ve been having..?

It actually has been a gorgeous evening, but unfortunately not for long enough. We woke up this morning to the tantalizing possibility of an EVA, however last night’s inclement weather didn’t subside soon enough and the ground was left rain-soaked and wind-battered, rendering it unstable for our entire EVA slot. If you can’t trust terra firma, who can you trust?

Instead, we spent the time grinding away on our research, adjusting to our new way of life, and oscillating in and out of latke-induced food comas. Kasey told us about her unsettling Stephen King novel. Vladimir taught us Durak, a traditional Russian card game that practically encourages over-the-top card-drawing theatrics. Cesare slaved over our Christmas Eve feast in the kitchen for hours, overwhelming the Habitat with some much needed pleasant aromas. Very much needed. We haven’t showered in 5 Sols.

I’ve been fasting since this morning so I can well and truly indulge in Cesare’s handiwork, but I think my stomach has begun to self-cannibalize. At a certain point, biological needs overpower literary ones. I’ve reached that point. Plus, what kind of crewmate would I be if I missed Christmas dinner?

And so, I bid you a wonderful Christmas Eve, and a jolly good pip pip cheerio!

Crew 236 Commander Report and Sol Summary 24-Dec-2021

[Commander Report, Sol Summary December 24th]

Crew 236 Sol Summary Report 24-DEC-2021

Sol: 5

Summary Title: Gotta Get Down on Friday

Author’s name: Kasey Hilton

Mission Status: Nominal

Sol Activity Summary: It’s Christmas Eve! We decided to sleep in a little bit this morning. Once everyone emerged from their cocoons, we feasted on a breakfast of latke (potato pancakes) cooked by our lovely journalist and GreenHab officer. There was an EVA scheduled for the afternoon, but due to the wind and rain, it was canceled. Instead, the afternoon was spent catching up on some reading. Crew 236 has quite the spread of reading materials here on Mars, let me tell ya! Tonight, we are having a holiday dinner consisting of homemade lasagna, garlic bread, zucchini “meatballs”, and beignets. Once we’ve demolished the spread, we plan to play card games and maybe watch a movie!

Look Ahead Plan: Tomorrow is Christmas Day! A quick EVA is planned for the afternoon and then we will be relaxing and celebrating together.

Anomalies in work: None

Weather: Rainy in the morning followed by a warm, windy, and overcast afternoon

Crew Physical Status: Feeling fun and festive

EVA: EVA was scheduled for this afternoon but canceled due to rain

Reports to be filed: Operations Report, GreenHab Report, EVA Report, EVA Request, Health and Safety Report, Sol Summary, and Journalist Report (with photos)

Support Requested: None

GreenHab Officer – December 24TH

[category: GreenHab-Report]

Crew 236 GreenHab Report 12242021
GreenHab Officer: Vladimir Zeltsman
Environmental Control: Around 17% Humidity
Average Temperatures: 70 degrees
Hours of Supplemental Light: 7 hours
Daily Water Usage For Crops: 9.5 gallons
Daily Water Usage For Research and/or Other Purposes: None
Water In Blue Tank: 203.5 Gallons
Times of watering for crops: 11:30 am
Changes to Crops: None
Narrative: None
Harvest: 2g Basil, 37g Tomato, 1g Green Onion
Support/supplies needed: None

GreenHab Officer – December 24TH

MJ!

Sorry to reply so late, we’ve been fixing a watery situation at the outpost.
I think you might have the wrong email to send the reports. Gimme a second

On Fri, Dec 24, 2021, 19:23 MJ Marggraff <mj.red.planet.rising> wrote:

[category: GreenHab-Report]

Crew 236 GreenHab Report 12242021
GreenHab Officer: Vladimir Zeltsman
Environmental Control: Around 17% Humidity
Average Temperatures: 70 degrees
Hours of Supplemental Light: 7 hours
Daily Water Usage For Crops: 9.5 gallons
Daily Water Usage For Research and/or Other Purposes: None
Water In Blue Tank: 203.5 Gallons
Times of watering for crops: 11:30 am
Changes to Crops: None
Narrative: None
Harvest: 2g Basil, 37g Tomato, 1g Green Onion
Support/supplies needed: None

Astronomy Report – December 24th

Astronomy Report

Name: Dylan Dilger

Crew: 236
Date: 12/24/2021

MDRS ROBOTIC OBSERVATORY

Robotic Telescope Requested (choose one) Montana Learning Center

Objects Viewed: M35, IC434 (Horsehead Nebula), M95, M96, NGC2244 (Rosette Nebula)

Problems Encountered: Observatory live view showed clear skies when observation requests were submitted and about 1/4 cloudy skies confined to the horizon 2 hours later. Looking at the sample JPG of all observations shows no stars and no detail, so clouds likely interfered with all observations. On Skynet, it said the weather was good, but I don’t think it’s very all-encompassing or sensitive.

MUSK OBSERVATORY

Solar Features Observed: None

Problems Encountered: None